Messages 0xx


01: LAMPOLY uniformity test (Benjamin Chui)
02: Re: LAMPOLY uniformity test (Jim McVittie)
03: more on etch rate uniformity (aaronp)
04: removal of 5mm edge bead using lithography (Sharleen Beckwith)
05: post HBr rinse (Sharleen Beckwith)
06: Re: post HBr rinse (Jim McVittie)
07: Re: post HBr rinse (Jim McVittie)
08: Black Silicon (Theresa Kramer)
09: More Contact people for LamPoly Problem (Jim McVittie)
10: Test (Bill Murray)
11: Dummy Wafer Change (Jim McVittie)
12: Nitride etch rates? (Nancy Latta)
13: No More Dummy Wafers to be Provided (Jim McVittie)
14: Out of HBr (Jim McVittie)
15: gas availability (Len Booth)
16: HBr gas (Dick Crane)
17: HBr (Dick Crane)
18: HBr (Ray Seymour)
19: freon 116 (Ray Seymour)
20: freon 116 (Ray Seymour)
21: CF4 and SF6 (Jim McVittie)
22: Plasma Cleaning (Jim McVittie)
23: Update on Lam Particle Problem (Jim McVittie)
24: my reservation from noon until 3 pm has been cancelled (Sharleen Beckwith)
25: New Log Sheet, RF Time, Wafer Count (Jim McVittie)
26: LAMpoly downtime (Len Booth)
27: LAMpoly downtime - update 04/12/01 (Len Booth)
28: Seeing High Oxide Etch Rates Need Old Rates (Jim McVittie)
29: Problems with Thick Resist in Etchers (Jim McVittie)
30: Status of Oxide etch rates in Lampoly (Jim McVittie)
31: Re: Status of Oxide etch rates in Lampoly (Aaron Partridge)
32: Particle problem resolved (Brian Joseph Greene)
33: LAMpoly shutdown reminder (Len Booth)
34: LAMpoly - update on particles, Friday 04/20 (Len Booth)
35: LAMpoly shutdown update 04/24/01 (Len Booth)
36: particle count update (Sharleen Beckwith)
37: Re: particle count update (Dan Grupp)
38: Toxic gas shutdown Aug. 6 (Dick Crane)
39: Toxic gas tests complete (Dick Crane)
40: HBr gas is not available at this time. (Len Booth)
41: HBr is available again (Len Booth)
42: HBR status (ray seymour)
43: Re: Shutdown lampoly 2002-01-16 10:10:52: won't pumpdown (Sharleen Beckwith)
44: Wet clean (Cesar Baxter)
45: LAMpoly status 04/02/02 (Len Booth)
46: LAMpoly status 04/08/02 (Len Booth)
47: LAMpoly update 04/12/02 (Len Booth)
48: gas bottle checks (Ray Seymour)
49: LAMpoly recipe files (Len Booth)
50: Re: LAMpoly recipe files (Jim McVittie)
51: LAMpoly status Sept 10,2002 (Len Booth)
52: LAMpoly status Sept 13, 2002 (Len Booth)
53: LAMpoly status Fri Oct 18 (Len Booth)
54: LAMpoly status Monday Oct 21 (Len Booth)
55: LAMpoly shutdown today (Len Booth)
56: Altered recipes (Benjamin Chui)
57: No Lam Etching Deeper Than 2 um (Jim McVittie)
58: Chlorine and BCl3 gas cylinder change (Elmer Enriquez)
59: Videotape (Mary Tang)
60: HBR leak- Poly etch not available. (Nancy Latta)
61: Status of HBr gas for poly etch (Dick Crane)
62: HBr gas status (Dick Crane)
63: HBr is back on-line (Dick Crane)
64: Lam is free from 3pm onward today (Robin King)
65: reservation removed friday afternoon (Ginel Hill)
66: reservation cancelled (Gladys)
67: HBr panel replacement (Elmer Enriquez)
68: Re: HBr panel replacement (Elmer Enriquez)
69: Update (Ceasar Baxter)
70: Eat to live, not live to eat (Gore)
71: Lampoly (Cesar M Baxter)
72: Update (Cesar M Baxter)
73: HBR problem (Cesar M Baxter)
74: Cl2 problem (Cesar M Baxter)
75: Cl2 problem (Cesar M Baxter)
76: ETA (Robin King)
77: Re: ETA (Dick Crane)
78: Shutdown (Cesar M Baxter)
79: Turbo Pump (Cesar M Baxter)
80: turbo update (Elmer Enriquez)
81: Lampoly update (Elmer Enriquez)
82: Lampoly update 6/21 (Elmer Enriquez)
83: Re: Lampoly update 6/21 (madan sanga)
84: Lampoly update (Elmer Enriquez)
85: Lampoly update 6/28 (Elmer Enriquez)
86: Lampoly update 7/3 (Elmer Enriquez)
87: Lampoly Update 7/13 (Elmer Enriquez)
88: lampoly running well (Raymond Woo)
89: Lam Recipes (Cesar M Baxter)
90: CL2 (cbax.stanford.edu)
91: HBR status (Ted Berg)
92: Cl2 and HBr (Mary Tang)
93: Etcher and Cl status (Ed Myers)
94: Cl2 unavailable later today (Mary Tang)
95: Chlorine problem (Mary Tang)
96: Cl2 and HBr update (Mary Tang)
97: Strange profiles reported... (Mary Tang)
98: No HBr (Ed Myers)
99: Re: No HBr (John Shott)